OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (anonymous):

well point-slope form is y-y1=m(x-x1) y1 is the y-coordinate x1 is the x-coordinate m is the slope y-___=___(x-____) fill in the blanks and see what you get

OpenStudy (anonymous):

would -3 be x2 and y2

OpenStudy (anonymous):

no -3 would go where the m is

OpenStudy (anonymous):

how do i set it up

OpenStudy (anonymous):

????

OpenStudy (anonymous):

just put 6 in for y1, -1 in for x1 and -3 in for m

OpenStudy (anonymous):

6-6=-3(-1-1

OpenStudy (anonymous):

close y-6=-3(x-(-1)) or y-6=-3(x+1) which is b

OpenStudy (anonymous):

thanks Spenser this helped me alot

OpenStudy (anonymous):

@mathman806 do u have a ps3

OpenStudy (anonymous):

no I wish I did though

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

yep